OpenStudy (anonymous):

Factor the following polynomial: x2 – 3x – 40

Parth (parthkohli):

Hint: 1 * -40 = -40 -8 * 5 = -40

OpenStudy (anonymous):

The answer please.

Parth (parthkohli):

Therefore, you get the following after splitting the middle term:\[x^2 - 8x + 5x - 40\]

Parth (parthkohli):

Factor by grouping.

OpenStudy (anonymous):

So is the answer (x – 8)(x + 5) ?

OpenStudy (baldymcgee6):

@ramirokool once again, is the function 2x or x^2

Parth (parthkohli):

Yeah, @ramirokool.

Parth (parthkohli):

@baldymcgee6 x^2.

OpenStudy (baldymcgee6):

\[f(x) = x^2– 3x – 40\] make sure when you post your function you post it correctly, it makes a huge difference.